Suppose that there is a monopoly cable TV company who offers two types of program: P1 and P2. There are two consumers (i = 1, 2) in this market, where consumer 1 has WTP of $12 for P1 and $10 for P2, while consumer 2 has WTP of $b for P1 and $14 for P2. The company must charge the same prices to both consumers, either for the two program types separately or for the bundle of both program types. Assume that there is no marginal cost of broadcasting TV program for the company.

Q8. When b = 1,

(a) setting prices for the two program types separately is superior and only consumer 2 will purchase P1 under the optimal pricing.

(b) setting a price for the bundle of both program types is superior and the bundle price will be $22 under the optimal pricing.

(c) setting a price for the bundle of both program types is superior and the bundle price will be $19 under the optimal pricing.

(d) setting prices for the two program types separately is superior and only consumer 1 will purchase P1 under the optimal pricing.

(e) setting prices for the two program types separately is superior and only consumer 2 will purchase P2 under the optimal pricing.

Q9. When b = 5,

(a) setting prices for the two program types separately is superior and only consumer 2 will purchase P1 under the optimal pricing.

(b) setting a price for the bundle of both program types is superior and the bundle price will be $22 under the optimal pricing.

(c) setting a price for the bundle of both program types is superior and the bundle price will be $19 under the optimal pricing.

(d) setting prices for the two program types separately is superior and only consumer 1 will purchase P1 under the optimal pricing.

(e) setting prices for the two program types separately is superior and only consumer 2 will purchase P2 under the optimal pricing.

Q10. When b = 10, 3

(a) setting prices for the two program types separately is superior and only consumer 2 will purchase P1 under the optimal pricing.

(b) setting a price for the bundle of both program types is superior and the bundle price will be $22 under the optimal pricing.

(c) setting a price for the bundle of both program types is superior and the bundle price will be $19 under the optimal pricing.

(d) setting prices for the two program types separately is superior and only consumer 1 will purchase P1 under the optimal pricing.

(e) setting prices for the two program types separately is superior and only consumer 2 will purchase P2 under the optimal pricing.
Answer #1

There are two programs P1 and P2. Consumer 1 has WTP = $12 for P1 and WTP for program 2 P2 = $10.WTP for the bundle by consumer 1 is 12 + 10 = $22 Consumer 2 has WTP = $b for P1 and WTP for program 2 P2 = $14. WTP for the bundle is ($14 + $b).

8) When b is 1, the WTP = $1 for P1 by consumer 2, the bundled price is $15 which will maximize profits. This will earn a profit of $30. Separate pricing will be $12 for P1 and $10 for P2 so that profits are now 12*1 + 10*2 = $32. Here separate pricing is superior. Consumer 2 has WTP = $1 for P1 and so it never buys P1. Correct option is A)

9) When b is 5, the WTP = $5 for P1 by consumer 2, the bundled price is $19 which will maximize profits. This will earn a profit of $38. Separate pricing will be $12 for P1 and $10 for P2 so that profits are now 12*1 + 10*2 = $32. Here separate pricing is inferior and bundled price of $19 is chosen. Option C is correct.

10) When b is 10, the WTP = $10 for P1 by consumer 2, the bundled price is $22 which will maximize profits. This will earn a profit of $44. Separate pricing will be $10 for P1 and $10 for P2 so that profits are now 10*2 + 10*2 = $40. Here separate pricing is inferior and bundled price of $19 is chosen. Option B is correct

Hope this helps you God bless you